OpenStudy (ashley1nonly):

An electron with initial velocity v0 = 1.5 ✕ 105m/s enters a region 1.0 cm long where it is electrically accelerated. It emerges with velocity v = 6.60 ✕ 106 m/s. What was its acceleration, assumed constant?

OpenStudy (ashley1nonly):

can i use the d=rt to get the time .01= 6.60*10^6 * t

OpenStudy (ashley1nonly):

t=1.5625*10^-9

OpenStudy (ashley1nonly):

can you help

jimthompson5910 (jim_thompson5910):

you can use a chart like this one http://www.studyphysics.ca/newnotes/20/unit01_kinematicsdynamics/chp04_acceleration/images/formulas.GIF to help you figure out what formula to use

OpenStudy (ashley1nonly):

i used the equatioon (v^2-v^0)/2x

jimthompson5910 (jim_thompson5910):

in this case, we don't need or care about the time t so look for the row that has an X under the t column (that's row 4)

jimthompson5910 (jim_thompson5910):

so yeah, you'll use \[\Large V_f^2 = V_i^2 + 2ad\] and solve for 'a'. It sounds like you are using the right formula

OpenStudy (ashley1nonly):

yes and i got (6.60*10^6)^2 - (1.5*10^5)^2 /2(.01)

OpenStudy (ashley1nonly):

my answer still came out wrong

OpenStudy (ashley1nonly):

it supposed to be in m/s^2

jimthompson5910 (jim_thompson5910):

ok one sec

OpenStudy (ashley1nonly):

ok

jimthompson5910 (jim_thompson5910):

I'm getting this roughly 2.176875*10^15 is that what you got?

OpenStudy (ashley1nonly):

i got 2.1*10^13

jimthompson5910 (jim_thompson5910):

oh so if you use 2 sig figs, then I'd round 2.176875*10^15 to 2.2*10^15

jimthompson5910 (jim_thompson5910):

and I wouldn't round until all the calculations are done

OpenStudy (ashley1nonly):

so did i have the original set up right

jimthompson5910 (jim_thompson5910):

yes you have the right equation

jimthompson5910 (jim_thompson5910):

and you plugged in the values correctly

OpenStudy (ashley1nonly):

thanks

jimthompson5910 (jim_thompson5910):

if it still doesn't accept that answer, then perhaps there's a typo or some computer glitch (making the computer really picky, idk)

OpenStudy (ashley1nonly):

it right

jimthompson5910 (jim_thompson5910):

ok great